OpenStudy (anonymous):

I just need someone to check my work. Use the rational root theorem to list all possible rational roots of the polynomial equation. x^3+7x2-10x-3 My answers are + or - 1 and + or - 3. Are these right?

OpenStudy (michele_laino):

what is your equation, please?

OpenStudy (anonymous):

The equation is the one on top. x^3+7x^2-10x-3

OpenStudy (michele_laino):

it is a polynomial only, I don't see the "=" sign

OpenStudy (anonymous):

Oh yeah that's it. We're just given the polynomial and have to find the zeros. I'm pretty sure I got the correct answer tho.

OpenStudy (anonymous):

Just wanna make sure lol :)

OpenStudy (michele_laino):

if I set x=1, I get this: 1^3+7*1^2-10*1-3=1+7-10-3=-5 which is not zero so x=1 can not be a zero of your polynomial

OpenStudy (anonymous):

It says to find all the possible zeros. So I'd have to use the rational root theorem and that p over q thing. So the factors of 3 over the factors of 1 and I get + or - 3 and + or - 1. Those would be the POSSIBLE zeros correct?

OpenStudy (michele_laino):

ok! that's right, your possible values of p/q are \[\pm 1, \quad \pm3\]

OpenStudy (anonymous):

Thanks so much! :) Did you get my message btw?

OpenStudy (michele_laino):

yes! please wait a moment

OpenStudy (anonymous):

Ok lol, thanks :)

OpenStudy (michele_laino):

for your first exercise, you have to check what possible values of p/q satisfy the condition below: \[P(p/q)=0\] where P is your polynomial

OpenStudy (anonymous):

Oh ok, yeah I did use that. There's one question that asks for all the possible values and one where it asks for all the zeros. Thanks tho for checking my work :)

OpenStudy (michele_laino):

ok! what is your next exercise?

OpenStudy (anonymous):

Ok well one of them is: 13) A polynomial equation with rational root coefficients has the roots \[2+\sqrt{4,}\], \[2\sqrt{5}\], find two additional roots. Show all work.

OpenStudy (michele_laino):

I'm trying...

OpenStudy (anonymous):

Alright, thanks so much I really appreciate it. The next question I'll post on a separate thread so I can give you more medals.

OpenStudy (anonymous):

Alright! Thank you :)

OpenStudy (michele_laino):

Sorry, I have made an error! I retry

OpenStudy (anonymous):

Oh that's fine :) It's ok.

OpenStudy (michele_laino):

we can say that your polynomial can be factored as below: \[P(x)=(x-2\sqrt{5})(x-2-\sqrt{4})*q(x)\] where q(x) is another polynomial whose degrre is >=2

OpenStudy (michele_laino):

opps..whose degree is...

OpenStudy (anonymous):

Ok... so then I just solve for x?

OpenStudy (michele_laino):

possible value of p/q are for example: \[\frac{ 2\sqrt{5} }{ 1 },\frac{ \sqrt{5} }{ 2 }, \frac{ -\sqrt{5} }{ -2 }\] and so on, but I can not infer that, for example sqrt(5)/2 is a rational root of your equation

OpenStudy (anonymous):

Oh ok, yeah that makes sense.

OpenStudy (anonymous):

Thanks tho! I'll just plug them in and check :) I really appreciate it :)

OpenStudy (michele_laino):

thanks! :) I continue to try